# Rourkela - 769008, India

### v1

@@ Abstract
We study the relationship between the standard or extended thermodynamic phase structure of various AdS black holes and the Lyapunov exponents associated with the null and time-like geodesics. We consider dyonic, Bardeen, Gauss-Bonnet, and Lorentz-symmetry breaking massive gravity black holes and calculate the Lyapunov exponents of massless and massive particles in unstable circular geodesics close to the black hole. We find that the thermal profile of the Lyapunov exponents exhibits distinct behaviour in the small and large black hole phases and can encompass certain aspects of the van der Waals type small/large black hole phase transition. We further analyse the properties of Lyapunov exponents as an order parameter and find that its critical exponent is $1/2$, near the critical point for all black holes considered here.
1 Introduction
@@
Chaos theory is a branch of physics and mathematics that explores the behaviour of dynamic systems highly sensitive to initial conditions, leading to ostensibly random and unpredictable outcomes. The theory originated from the pioneering work of mathematician and meteorologist Edward Lorenz Lorenz:1963yb ; Lorenzelli2014Apr , who discovered that small changes in the initial conditions could lead to vastly different outcomes in non-integrable systems, giving rise to the famous term “butterfly effect”. Central to chaos theory is the concept of the Lyapunov exponent Lyapunov1992Mar , which measures the rate of divergence or convergence of nearby trajectories in a dynamical system sandri1996 . Positive Lyapunov exponents indicate divergence, suggesting sensitive dependence on initial conditions, which is synonymous with chaos, while negative exponents imply stability and convergence. Lyapunov exponents provide a quantitative measure of predictability in a system; higher values signify greater unpredictability and chaotic dynamics. This mathematical tool has applications across various scientific disciplines, including physics, biology, economics, and meteorology, helping researchers understand and model the inherent complexity of natural phenomena Colangelo:2021kmn ; Shukla:2023pbp ; ferriere1995lyapunov ; dechert1992lyapunov ; vannitsem2017predictability .
@@
There has been a significant interest in investigating phase transitions using chaos diagnosis tools in recent years. This investigation extends to both classical and quantum domains and has attracted great interest in many areas of physics, such as condensed matter and quantum information physics. The prime examples include the interplay of chaos and the quantum phase transition in the Dicke model Emary:2003zza , Sachdev-Ye-Kitaev model (SYK) and its variants Sorokhaibam:2019qho ; Davis:2022iqi , long-range models of coupled oscillators Miritello:2008zd , finite Fermi and quantum dot systems Heiss:1991zza ; emary2003chaos , etc. Two diagnostic tools for quantum chaos, Out-of-Time Order Correlators (OTOCs) and Krylov complexity have also gathered prominence as subjects of current scholarly interest daug2019detection ; sun2020out ; wang2019probing ; huh2021diagnosing ; anegawa2024krylov .
@@
Lately, this correlation of chaos and phase transitions has also left its footprints in general relativity and black hole physics. The bonafide table-top laboratories that investigate such correlation between chaos and phase transition are the black holes in AdS (Anti-de Sitter) spaces. In particular, black holes in AdS spaces are not only usually thermodynamically stable but also undergo interesting phase transitions, such as Hawking/Page or small/large black hole phase transitions, and exhibit a rich phase structure hawking1983thermodynamics ; chamblin1999holography . The latter transition, in particular, is of great interest as it generally involves a first-order transition line terminating at a second-order critical point, with critical exponents identified with the mean-field type, thereby emulating the standard van der Waals type phase transition of the liquid-gas system in black hole context. These interesting phase transitions are not typically associated and observed with their asymptotically flat counterparts. After the initial finding of chamblin1999holography on charged RN-AdS black hole, such small/large phase transitions have been observed in various AdS gravity models in different contexts. The examples are higher curvature Gauss-Bonnet black holes Cai:2013qga ; Dolan:2014vba , dyonic black holes lu2013ads ; Hartnoll:2007ai ; rasheed1995rotating ; tseytlin1996extreme ; cheng1993dyonic ; priyadarshinee2021analytic ; chaturvedi2017thermodynamic ; dutta2013dyonic , nonlinear electrodynamics based regular Bardeen black holes Balart:2014cga , $R$-charged gauge supergravity inspired STU black holes Behrndt:1998jd ; Duff:1999gh ; Cvetic:1999xp ; Sahay:2010yq , hairy lower-dimensional black holes Priyadarshinee:2023cmi , quasi-topological gravity Hennigar:2015esa , $f(R)$ gravity Chen:2013ce , etc. The phase structure of AdS black holes becomes even more interesting in the context of extended phase space thermodynamics, where the cosmological constant is interpreted as thermodynamic pressure and treated as another thermodynamic variable Kubiznak:2012wp ; Dolan:2014jva ; Kastor:2009wy ; Dolan:2011xt ; Dolan:2012jh . See Kubiznak:2016qmn ; Altamirano:2014tva for reviews on this interesting topic.
@@
Free energies are the standard approach to analysing the black hole phase transitions. Lately, there have been attempts to probe the black hole phase transition with indirect methods, like quasinormal modes (QNMs) Liu:2014gvf ; Shen:2007xk ; Rao:2007zzb ; He:2010zb ; Mahapatra:2016dae ; Priyadarshinee:2023exb , black hole shadow radius Belhaj:2020nqy ; Cai:2021uov ; Zhang:2019glo , photon orbits Wei:2017mwc ; Chabab:2019kfs ; Han:2018ooi , entanglement entropy Johnson:2013dka ; Dey:2015ytd ; Dudal:2018ztm ; Mahapatra:2019uql ; Jain:2022hxl ; Jain:2020rbb etc. In such indirect methods, the physical quantities exhibit both multi-valuedness and discontinuity at the phase transition point, analogous to the thermodynamic variables in ordinary thermodynamics. Lyapunov exponents are another step in searching for indirect black hole phase transitions probes. Indeed, the fact that the Lyapunov exponents of unstable null geodesics are closely related to the imaginary part of QNMs Cardoso:2008bp ; Guo:2021enm and that QNMs do exhibit noticeable changes near the phase transition indicates that pronounced changes may occur in the structure of Lyapunov exponents near the black hole phase transition.
@@
Figure 1: This shows a timeslice of the unstable circular geodesics of the massive particles designated by $\bullet$, $\bullet$ and $\bullet$ colours (with slightly perturbed specific initial positions and velocities) around a Schwarzschild black hole (with mass $M=0.5$ in natural units). The black sphere in the centre represents the black hole.
@@
Intriguing investigations have recently started on the correlation of black hole phase transition and Lyapunov exponent guo2022probing ; yang2023lyapunov ; lyu2023probing ; kumara2024lyapunov . It has been observed that the structure of black hole phase transitions is imprinted on the Lyapunov exponents of massive and massless probe particles moving around the black hole backgrounds in unstable circular geodesics, as illustrated in Fig. 1. In particular, the Lyapunov exponent exhibits multi-valuedness as a function of temperature, a scenario usually associated with the phase transition, and its thermal profile simulates the thermal profile of black hole entropy. Interestingly, the Lyapunov exponent also becomes single-valued at some critical model-dependent parameter value, thereby indirectly encoding the information about the second-order critical point.
@@
This is fascinating, considering that in the context of gauge/gravity duality, the small/ large black hole phase transition in AdS spaces corresponds to some large $N$ phase transition, $N$ being the number of colours, such as confined/deconfined phase transition, in the dual boundary field theory at strong couplings. [^1] The small black hole phase does not correspond to the confined phase; however, it approximately shares many of its features. For more discussion on this, see Dudal:2017max . Therefore, the behaviour of the Lyapunov exponent provides another way to look at the structure of confined/deconfined phases at strong couplings. This, in turn, may also be useful for studying chaotic properties of strongly coupled QCD systems, an issue not directly accessible via standard computations like lattice QCD.
@@
In the present work, we further test this interplay of black hole phase transition and Lyapunov exponents in various AdS black holes. Specifically, we concentrate on dyonic black holes, Bardeen black holes, Gauss-Bonnet black holes, and massive gravity black holes. These choices of black holes are motivated by various physical reasons. For instance, the phase structure of dyonic black holes, which contain additional magnetic charge, is much richer than RN-AdS black holes lu2013ads ; Hartnoll:2007ai ; rasheed1995rotating ; tseytlin1996extreme ; cheng1993dyonic ; priyadarshinee2021analytic ; chaturvedi2017thermodynamic ; dutta2013dyonic . Hence, it would be interesting to see how the magnetic charge modifies the thermal profile of the Lyapunov exponent. Similarly, the everywhere regular Bardeen black holes are sourced by non-linear electrodynamics and do not contain any curvature singularity Ayon-Beato:2000mjt . Therefore, it would be interesting to see how the Lyapunov exponent behaves near the phase transition involving regular black holes. On the other hand, the Gauss-Bonnet black holes are the simplest prototype of higher-order curvature gravity theories Boulware:1985wk ; Deser:2002jk . Such higher curvature theories are expected to play an essential role in quantum gravity and arise naturally from the low-energy expansion of string theory Gross:1986mw . The Gauss-Bonnet black holes also exhibit a van der Waals-type phase transition, and it is certainly interesting to investigate whether the Lyapunov exponent can probe such a phase transition in higher curvature gravity theories or not Cai:2013qga ; Dolan:2014vba ; lyu2023probing . Similarly, massive gravity theory corresponds to a non-trivial modification of general relativity in which the graviton acquires a mass. This modification changes general relativity by weakening it at large scales. Still, it leads to the exact prediction as general relativity at small scales Fierz:1939ix ; deRham:2010ik ; deRham:2010kj ; Dvali:2000hr ; Bergshoeff:2009hq ; deRham:2014zqa ; Hinterbichler:2011tt . The thermodynamics of AdS massive gravity further reveals novel phase transitions in this system Cai:2014znn ; Xu:2015rfa ; Hendi:2017fxp ; Zou:2016sab ; Hendi:2016vux ; Fernando:2016sps .
@@
In all the above-mentioned black holes, a critical value of a parameter (usually the black hole charge) appears below, and these black holes exhibit the small/large black hole phase transition. However, above this critical value, no such phase transition occurs. This can be true with and without extended thermodynamics and in different ensembles. We compute the Lyapunov exponent associated with both massive and massless particles propagating in these black hole backgrounds and find that, in all cases, the thermal profile of the Lyapunov exponent captures the essence of the phase transition. In particular, the thermal profile of the Lyapunov exponent is multi-valued for smaller parameter values, whereas it is single-valued for larger parameter values. Interestingly, this structure of the Lyapunov exponent also allows us to pinpoint the critical value of the parameter, thereby indirectly providing information about the second-order critical point. Following lyu2023probing , we further define the Lyapunov exponent and compute the corresponding critical exponent at the phase transition point. This critical exponent has a value of one-half in all black hole backgrounds considered here. This is true for both standard and extended-phase space thermodynamics.
@@
The paper is organised as follows. In Sec. 2, we discuss the methodology and main recipe for computing the Lyapunov exponents associated with massless and massive particles. We then discuss the black hole phase transition and its interplay with the Lyapunov exponent (both for massless and massive particles) for various charged black holes in Sec. 3. The use of discontinuity in the Lyapunov exponent as an order parameter is illustrated in Sec. 4. Finally, we conclude our paper by highlighting the principal results and scope of future work in Sec. 5.
2 Methodology
@@
In this section, we briefly establish the calculation relating the principal Lyapunov exponent $\lambda$ of massless and massive particles to the effective potential of unstable orbits. We will mainly follow the algorithm suggested in Cardoso:2008bp ; guo2022probing , and more details can be found there. For this purpose, we start with the following spherically symmetric static AdS black hole background
@@
We follow the procedure used by Chandrasekhar chandrasekhar1991selected to calculate the Lagrangian of the particle’s geodesic motion for the above spacetime (1),
@@
Note that $\epsilon=-1$, resp. 00 corresponds to massive, resp. massless particles moving along time-like, resp. null geodesics. Now, restricting our attention to the equatorial plane $\theta=\phi=\pi/2$, we have
@@
where $K$ is the linear stability matrix Cornish:2003ig and is given by,
@@
The relation which gives the principal Lyapunov exponent for the circular orbits is then derived as Cardoso:2008bp
@@
3 Thermodynamics and Phase Structure of charged black holes based on Lyapunov exponents
@@
3.1 Dyonic Black Holes
Because of electromagnetic duality in four dimensions, it is feasible to obtain black hole solutions with electric and magnetic charges. Such black holes are called dyonic black holes and can be considered as the simplest generalisations of RN-AdS black holes lu2013ads ; rasheed1995rotating ; tseytlin1996extreme ; cheng1993dyonic ; priyadarshinee2021analytic . The magnetic charge not only enriches the thermodynamic phase diagram of RN-AdS black holes but also acquaints the dual boundary theory with a background magnetic field, thereby making these solutions relevant for the holographic study of, for instance, the Hall effect Hartnoll:2007ai , ferromagnetism dutta2013dyonic etc.
Dyonic black holes are solutions of the Einstein-Maxwell gravity system. The details of the solutions can be found in dutta2013dyonic . Its line element reads as
@@ 3.1 Dyonic Black Holes
Below, we discuss the standard thermodynamics of the dyonic black hole in the grand-canonical ensemble. One can perform a similar analysis in the extended phase space as well. Since the thermodynamic results of the extended phase space are quite similar to the above-discussed results, we, therefore, concentrate only on normal thermodynamics without losing any generality.
@@
The curve projected onto the $V_{\text{eff}}(r)=-1500$ plane represents the extrema of the effective potential for the massive particle ($\epsilon=-1$). The reddish yellow part represents the maxima or the radii of the unstable time-like circular geodesics given by $V^{{}^{\prime}}_{\text{eff}}(r_{0})=0$ and $V^{{}^{\prime\prime}}_{\text{eff}}(r_{0})<0$, while the blueish-white part represents the minima or the radii of the stable time like circular geodesic given by $V^{{}^{\prime}}_{\text{eff}}(r_{0})=0$ and $V^{{}^{\prime\prime}}_{\text{eff}}(r_{0})>0$. As we increase the $r_{h}$ slowly, the maxima and the minima shift. At a particular $r_{h}=0.827$, they become equal, represented by the black line on the curve, after which there are no extrema for the massive particle. We are only interested in the unstable circular geodesics, which are important in calculating the Lyapunov exponents.
3.1.1 Massless particles

### v2

# Interplay between the Lyapunov exponents and phase transitions of charged AdS black holes

@@ Abstract
We study the relationship between the standard or extended thermodynamic phase structure of various AdS black holes and the Lyapunov exponents associated with the null and timelike geodesics. We consider dyonic, Bardeen, Gauss-Bonnet, and Lorentz-symmetry breaking massive gravity black holes and calculate the Lyapunov exponents of massless and massive particles in unstable circular geodesics close to the black hole. We find that the thermal profile of the Lyapunov exponents exhibits distinct behaviour in the small and large black hole phases and can encompass certain aspects of the van der Waals type small/large black hole phase transition. We further analyse the properties of Lyapunov exponents as an order parameter and find that its critical exponent is 1/2, near the critical point for all black holes considered here.
I Introduction
@@
Chaos theory is a branch of physics and mathematics that explores the behaviour of dynamic systems highly sensitive to initial conditions, leading to ostensibly random and unpredictable outcomes. The theory originated from the pioneering work of mathematician and meteorologist Edward Lorenz [1, 2], who discovered that small changes in the initial conditions could lead to vastly different outcomes in non-integrable systems, giving rise to the famous term “butterfly effect”. It has applications in many scientific fields and is also extensively used in the nascent field of holography [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. For a summary of recent developments in holographic chaos, let us refer to the review paper [16]. Central to chaos theory is the concept of the Lyapunov exponent [17], which measures the rate of divergence or convergence of nearby trajectories in a dynamical system [18]. Positive Lyapunov exponents indicate divergence, suggesting sensitive dependence on initial conditions, which is synonymous with chaos, while negative exponents imply stability and convergence. Lyapunov exponents provide a quantitative measure of predictability in a system; higher values signify greater unpredictability and chaotic dynamics. This mathematical tool has applications across various scientific disciplines, including physics, biology, economics, and meteorology, helping researchers understand and model the inherent complexity of natural phenomena [19, 20, 21, 22, 23].
@@
There has been a significant interest in investigating phase transitions using chaos diagnosis tools in recent years. This investigation extends to both classical and quantum domains and has attracted great interest in many areas of physics, such as condensed matter and quantum information physics. The prime examples include the interplay of chaos and the quantum phase transition in the Dicke model [24], Sachdev-Ye-Kitaev model (SYK) and its variants [25, 26], long-range models of coupled oscillators [27], finite Fermi and quantum dot systems [28, 29], etc. Two diagnostic tools for quantum chaos, Out-of-Time Order Correlators (OTOCs) and Krylov complexity have also gathered prominence as subjects of current scholarly interest [30, 31, 32, 33, 34].
@@
Lately, this correlation of chaos and phase transitions has also left its footprints in general relativity and black hole physics. The bonafide table-top laboratories that investigate such correlation between chaos and phase transition are the black holes in AdS (Anti-de Sitter) spaces. In particular, black holes in AdS spaces are not only usually thermodynamically stable but also undergo interesting phase transitions, such as Hawking/Page or small/large black hole phase transitions, and exhibit a rich phase structure [35, 36]. The latter transition, in particular, is of great interest as it generally involves a first-order transition line terminating at a second-order critical point, with critical exponents identified with the mean-field type, thereby emulating the standard van der Waals type phase transition of the liquid-gas system in black hole context. These interesting phase transitions are not typically associated and observed with their asymptotically flat counterparts. After the initial finding of [36] on charged RN-AdS black hole, such small/large phase transitions have been observed in various AdS gravity models in different contexts. The examples are higher curvature Gauss-Bonnet black holes [37, 38], dyonic black holes [39, 40, 41, 42, 43, 44, 45, 46], nonlinear electrodynamics based regular Bardeen black holes [47], $R$-charged gauge supergravity inspired STU black holes [48, 49, 50, 51], hairy lower-dimensional black holes [52], quasi-topological gravity [53], $f(R)$ gravity [54], etc. The phase structure of AdS black holes becomes even more interesting in the context of extended phase space thermodynamics, where the cosmological constant is interpreted as thermodynamic pressure and treated as another thermodynamic variable [55, 56, 57, 58, 59]. See [60, 61] for reviews on this interesting topic.
@@
Free energies are the standard approach to analysing the black hole phase transitions. Lately, researchers have actively sought to explore phase transitions of black holes through indirect methods, see for instance [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79] etc. In such indirect methods, the physical quantities exhibit both multi-valuedness and discontinuity at the phase transition point, analogous to the thermodynamic variables in ordinary thermodynamics. Lyapunov exponents are another step in searching for indirect black hole phase transition probes. Indeed, the fact that recent studies have actively demonstrated a close relationship between the imaginary part of quasinormal modes (QNMs) and the Lyapunov exponents of unstable null geodesics [80, 81] and that QNMs do exhibit noticeable changes near the phase transition indicates that pronounced modifications may occur in the structure of Lyapunov exponents near the black hole phase transition.
@@
Figure 1: This shows a time-slice of the unstable circular geodesics of the massive particles designated by $\bullet$, $\bullet$ and $\bullet$ colours (with slightly perturbed specific initial positions and velocities) around a Schwarzschild black hole (with mass $M=0.5$ in natural units). The black sphere in the centre represents the black hole.
@@
Intriguing investigations have recently started on the correlation of black hole phase transition and Lyapunov exponent [82, 83, 84, 85, 86, 87]. It has been observed that the structure of black hole phase transitions is imprinted on the Lyapunov exponents of massive and massless probe particles moving around the black hole backgrounds in unstable circular geodesics. For example, we have shown a timeslice of unstable circular geodesics of massive particles around a Schwarzschild black hole in Fig. 1. In particular, the Lyapunov exponent exhibits multi-valuedness as a function of temperature, a scenario usually associated with the phase transition, and its thermal profile simulates the thermal profile of black hole entropy. Interestingly, the Lyapunov exponent also becomes single-valued at some critical model-dependent parameter value, thereby indirectly encoding the information about the second-order critical point.
@@
This is fascinating, considering that in the context of gauge/gravity duality, the small/ large black hole phase transition in AdS spaces corresponds to some large $N$ phase transition, $N$ being the number of colours, such as confined/deconfined phase transition, in the dual boundary field theory at strong couplings. [^1] The small black hole phase does not correspond to the confined phase; however, it approximately shares many of its features. For more discussion on this, see [88]. Therefore, the behaviour of the Lyapunov exponent provides another way to look at the structure of confined/deconfined phases at strong couplings. This, in turn, may also be useful for studying chaotic properties of strongly coupled QCD systems, an issue not directly accessible via standard computations like lattice QCD.
@@
In the present work, we further test this interplay of black hole phase transition and Lyapunov exponents in various AdS black holes. Specifically, we concentrate on dyonic black holes, Bardeen black holes, Gauss-Bonnet black holes, and massive gravity black holes. These choices of black holes are motivated by various physical reasons. For instance, the phase structure of dyonic black holes, which contain additional magnetic charge, is much richer than RN-AdS black holes [39, 40, 41, 42, 43, 44, 45, 46]. Hence, it would be interesting to see how the magnetic charge modifies the thermal profile of the Lyapunov exponent. Similarly, the everywhere regular Bardeen black holes are sourced by non-linear electrodynamics and do not contain any curvature singularity [89]. Therefore, it would be interesting to see how the Lyapunov exponent behaves near the phase transition involving regular black holes. On the other hand, the Gauss-Bonnet black holes are the simplest prototype of higher-order curvature gravity theories [90, 91]. Such higher curvature theories are expected to play an essential role in quantum gravity and arise naturally from the low-energy expansion of string theory [92]. The Gauss-Bonnet black holes also exhibit a van der Waals-type phase transition, and it is undoubtedly interesting to investigate whether the Lyapunov exponent can probe such a phase transition in higher curvature gravity theories or not [37, 38, 84]. Similarly, massive gravity theory corresponds to a non-trivial modification of general relativity in which the graviton acquires a mass. This modification changes general relativity by weakening it at large scales. Still, it leads to the exact prediction as general relativity at small scales [93, 94, 95, 96, 97, 98, 99]. The thermodynamics of AdS massive gravity further reveals novel phase transitions in this system [100, 101, 102, 103, 104, 105].
@@
In all the above-mentioned black holes, a critical value of a parameter (usually the black hole charge) appears below, and these black holes exhibit the small/large black hole phase transition. However, above this critical value, no such phase transition occurs. This can be true with and without extended thermodynamics and in different ensembles. We compute the Lyapunov exponent associated with both massive and massless particles propagating in these black hole backgrounds and find that, in all cases, the thermal profile of the Lyapunov exponent captures the essence of the phase transition. In particular, the thermal profile of the Lyapunov exponent is multi-valued for smaller parameter values, whereas it is single-valued for larger parameter values. Interestingly, this structure of the Lyapunov exponent also allows us to pinpoint the critical value of the parameter, thereby indirectly providing information about the second-order critical point. Following [84], we further define the Lyapunov exponent and compute the corresponding critical exponent at the phase transition point. This critical exponent has a value of one-half in all black hole backgrounds considered here. This is true for both standard and extended-phase space thermodynamics.
@@
The paper is organised as follows. In Sec. II, we discuss the methodology and main recipe for computing the Lyapunov exponents associated with massless and massive particles. We then discuss the black hole phase transition and its interplay with the Lyapunov exponent (both for massless and massive particles) for various charged black holes in Sec. III. The use of discontinuity in the Lyapunov exponent as an order parameter is illustrated in Sec. IV. Finally, we conclude our paper by highlighting the principal results in Sec. V.
II Methodology
@@
In this section, we briefly establish the calculation relating the principal Lyapunov exponent $\lambda$ of massless and massive particles to the effective potential of unstable orbits. We will mainly follow the algorithm suggested in [80, 82], and more details can be found there. For this purpose, we start with the following spherically symmetric static AdS black hole background
@@
We follow the procedure used by Chandrasekhar [106] to calculate the Lagrangian of the particle’s geodesic motion for the above spacetime (1),
@@
Note that $\epsilon=-1$, resp. 00 corresponds to massive, resp. massless particles moving along timelike, resp. null geodesics. Now, restricting our attention to the equatorial plane $\theta=\phi=\pi/2$, we have
@@
where $K$ is the linear stability matrix [107] and is given by,
@@
The relation which gives the principal Lyapunov exponent for the circular orbits is then derived as [80]
@@
III Thermodynamics and Phase Structure of charged black holes based on Lyapunov exponents
@@
III.1 Dyonic Black Holes
Because of electromagnetic duality in four dimensions, it is feasible to obtain black hole solutions with electric and magnetic charges. Such black holes are called dyonic black holes and can be considered as the simplest generalisations of RN-AdS black holes [39, 41, 42, 43, 44]. The magnetic charge not only enriches the thermodynamic phase diagram of RN-AdS black holes but also acquaints the dual boundary theory with a background magnetic field, thereby making these solutions relevant for the holographic study of, for instance, the Hall effect [40], ferromagnetism [46] etc.
Dyonic black holes are solutions of the Einstein-Maxwell gravity system. The details of the solutions can be found in [46]. Its line element reads as
@@ III.1 Dyonic Black Holes
Below, we discuss the standard thermodynamics of the dyonic black hole in the grand-canonical ensemble. One can perform a similar analysis in the extended phase space as well. Since the thermodynamic results of the extended phase space are quite identical to the above-discussed results, we, therefore, concentrate only on standard thermodynamics without losing any generality.
@@
The curve projected onto the $V_{\text{eff}}(r)=-1500$ plane represents the extrema of the effective potential for the massive particle ($\epsilon=-1$). The reddish yellow part represents the maxima or the radii of the unstable timelike circular geodesics given by $V^{{}^{\prime}}_{\text{eff}}(r_{0})=0$ and $V^{{}^{\prime\prime}}_{\text{eff}}(r_{0})<0$, while the blueish-white part represents the minima or the radii of the stable time like circular geodesic given by $V^{{}^{\prime}}_{\text{eff}}(r_{0})=0$ and $V^{{}^{\prime\prime}}_{\text{eff}}(r_{0})>0$. As we increase the $r_{h}$ slowly, the maxima and the minima shift. At a particular $r_{h}=0.827$, they become equal, represented by the black line on the curve, after which there are no extrema for the massive particle. We are only interested in the unstable circular geodesics, which are important in calculating the Lyapunov exponents.
III.1.1 Massless particles